THe Kb of ammonia is 1.76*10-5. What is the pH of a buffer which is prepared by combining 50.0 mL of 1.21 M ammonia and 45.0 mL of 0.86 M ammonium nitrate?

A) 9.44

B) 9.05

C) 9.25

D) None of the above

E) 4.94
Answer #1

Henderson- Hanelbach equation pOH = pkz + log [NHG] INH₂] Here [NH] = 0.86 mol x 45 ml 45 mlt som [NHU J = 0.86*45 95 [NH3] =